OpenStudy (anonymous):

f(x)=1/x and g(x)=x^2-4x two numbers are not in the domain of f°g

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(x^2-4x)=\frac{1}{x^2-4x}\]

OpenStudy (anonymous):

since you cannot divide by zero, the two numbers you are looking for are the solutions to \[x^2-4x=0\]

OpenStudy (anonymous):

@satellite73 that is wrong

OpenStudy (anonymous):

what is wrong?

OpenStudy (anonymous):

the answer

OpenStudy (anonymous):

what answer?

OpenStudy (anonymous):

idk im confused

OpenStudy (anonymous):

can you help @satellite73

OpenStudy (anonymous):

i found out its (4,0)

OpenStudy (anonymous):

@satellite73 thank you